# Bias in physics peer recognition does not explain gaps in perceived recognition

### v2

@@ Abstract
Gaining recognition as a physics person from peers is an important contributor to undergraduate students’ physics identity and, in turn, their success in physics courses. Previous research has separately demonstrated that women perceive less recognition from peers than men in their physics courses (perceived recognition) and that women receive fewer nominations from their peers as strong in their physics course than men (received recognition). The relationship between perceived and received peer recognition, however, is not well understood. Here we test three plausible models for this relationship. We conduct a large-scale, quantitative study of over 1,600 students enrolled in introductory physics courses at eight different institutions. We directly compare student gender, perceived recognition, and received recognition, controlling for other student demographics and course-level variability. Results show with high precision that, for students receiving the same amount of recognition, and having the same race or ethnicity, academic year, and major, women report significantly lower perceived recognition than men. These findings offer important implications for the design of effective instructional interventions.
The under-representation of women and people of color in undergraduate physics courses and, subsequently, in physics careers is well-documented [1, 2, 3, 4]. Broadening the population of physicists would give voice to new and innovative ideas that are critical to scientific advancements [5, 6, 7]. While many structural barriers to entry and retention in physics exist for women in particular, one significant challenge is that women may both perceive [8, 9, 10, 11, 12] and receive [13, 14, 15, 16, 17] less recognition from their physics peers about their abilities as physicists than men. Such recognition from peers directly correlates with students’ development of their physics identity (i.e., the extent to which they believe they are a “physics person” [18]), and, subsequently, their persistence in physics courses and careers [19, 9, 10, 8, 20].
Existing research, however, has only examined gender differences in students’ perceived and received peer recognition separately [13, 14, 15, 16, 17, 8, 9, 10, 11, 12]. Understanding the nuance of how students internalize the peer recognition they actually receive is crucial for identifying effective instructional interventions aimed at mitigating recognition gaps between men and women. In this study, therefore, we directly compare the extent to which students feel recognized by their peers (perceived recognition) to the number of nominations they receive from their peers for being strong in their physics course (which we assume, for now, is a proxy for received recognition), as indicated on surveys.
@@
We perform our analysis with 1,688 unique students across 27 introductory physics contexts (hereon referred to as individual “courses”): 15 instructional lecture courses (containing 1,039 total students, 559 men and 480 women) and 12 instructional lab courses (containing 1,652 total students, 794 men and 858 women; see Table 1). The courses come from eight PhD-granting institutions in the United States, including one Historically Black College or University and one Hispanic-Serving Institution. Four of the institutions are public and four are private. All 27 courses had a survey response rate above 70% of enrolled students. We exclude data collected from an additional 47 introductory physics courses (34 instructional lecture courses and 13 instructional lab courses, all from the same set of eight institutions) with a response rate below 70% because response rates below this threshold render our measure of received recognition too sensitive to missing data [21]. We also only include in the analysis students who self-identify as a man or a woman in order to make quantitative comparisons between these two groups.
We note that we examine students in physics courses pursuing a variety of majors and that such students may have different motivations for taking physics. Physics identity, however, is associated with students’ success in their physics courses, and thus is valuable for any student who is required to take physics. Moreover, physics identity has been shown to be important for the identities and career choices of engineering students, who comprise a large fraction of our sample [22]. Non-physics students’ experiences in physics courses can also shape their other disciplinary identities and intentions to pursue their chosen field of study [23].
@@ DISCUSSION
Yet another possibility is that men and women internalize recognition from their peers differently: either men over-estimate their peer recognition or women under-estimate their peer recognition, or some combination of the two. This may occur, for example, if men and women have biased self-assessments that then lead to biases in their perceptions of how their peers view them. Previous research has demonstrated such biases in students’ self-assessments, finding that women who earn high grades in their physics classes report similar levels of self-efficacy (i.e., beliefs about their capabilities to successfully carry out physics tasks) as men who earn much lower grades [31].
@@
These and similar lines of research will further illuminate how students internalize their peer recognition and why men and women internalize their peer recognition differently, as we observe in the large-scale analysis presented here. With these mechanisms, the research community can then design, implement, and evaluate instructional interventions that are effective in mitigating gender differences in perceived peer recognition. Such interventions will help more women and students from marginalized groups develop their physics identity, and, subsequently, better succeed in their physics courses.

### v3

@@ Abstract
Gaining recognition as a physics person from peers is an important contributor to undergraduate students’ physics identity and their success in physics courses. Previous research has separately demonstrated that women perceive less recognition from peers than men in their physics courses (perceived recognition) and that women receive fewer nominations from their peers as strong in their physics course than men (received recognition). The relationship between perceived and received peer recognition between men and women, however, is not well understood. Here we test three plausible models for this relationship. We conduct a large-scale, quantitative study of over 1,600 students enrolled in introductory physics courses at eight different institutions. We directly compare student gender, perceived recognition, and received recognition, controlling for other student demographics and course-level variability. Results show with high precision that, for students receiving the same amount of recognition, and having the same race or ethnicity, academic year, and major, women report significantly lower perceived recognition than men. These findings offer important implications for the design of effective instructional interventions.
The under-representation of women and people of color in undergraduate physics courses and, subsequently, in physics careers is well-documented [1, 2, 3, 4]. Broadening the population of physicists would give voice to new and innovative ideas that are critical to scientific advancements [5, 6, 7]. While many structural barriers to entry and retention in physics exist for women in particular, one significant challenge is that women may both perceive [8, 9, 10, 11, 12] and receive [13, 14, 15, 16, 17] less recognition from their physics peers about their abilities as physicists than men. Such recognition from peers directly correlates with students’ development of their physics identity (i.e., the extent to which they believe they are a “physics person” [18]) and their persistence in physics courses and careers [19, 9, 10, 8, 20].
Existing research in physics education, however, has only examined gender differences in students’ perceived and received peer recognition separately [13, 14, 15, 16, 17, 8, 9, 10, 11, 12]. Understanding the nuance of how students internalize the peer recognition they actually receive is crucial for identifying effective instructional interventions aimed at mitigating recognition gaps between men and women. In this study, therefore, we directly compare the extent to which students feel recognized by their peers (perceived recognition) to the number of nominations they receive from their peers for being strong in their physics course (which we assume is a proxy for received recognition), as indicated on surveys.
@@
We perform our analysis with 1,688 unique students across 27 introductory physics contexts (hereon referred to as individual “courses”): 15 instructional lecture courses (containing 1,039 total students, 559 men and 480 women) and 12 instructional lab courses (containing 1,652 total students, 794 men and 858 women; see Table 1). The courses come from eight PhD-granting institutions in the United States, including one Historically Black College or University and one Hispanic-Serving Institution. Four of the institutions are public and four are private. All 27 courses had a survey response rate above 70% of enrolled students. We exclude data collected from an additional 47 introductory physics courses (34 instructional lecture courses and 13 instructional lab courses, all from the same set of eight institutions) with a response rate below 70% because response rates below this threshold render our measure of received recognition too sensitive to missing data [21]. We also only include in the analysis students who self-identify as a man or a woman because prior work predicts recognition differences between these two groups–there are not yet any established expectations of recognition differences between other gender identities.
We note that we examine students in physics courses pursuing a variety of majors and that such students may have different motivations for taking physics. Physics identity, however, is associated with students’ success in their physics courses, and thus is valuable for any student who is required to take physics. Moreover, physics identity has been shown to be important for the career choices of engineering students, who comprise a large fraction of our sample [22]. Non-physics students’ experiences in physics courses can also shape their other disciplinary identities and intentions to pursue their chosen field of study [23].
@@ DISCUSSION
Yet another possibility is that men and women internalize recognition from their peers differently: either men over-estimate their peer recognition or women under-estimate their peer recognition, or a combination of the two. This phenomenon may be a form of the Dunning-Kruger effect [31] where men and women have biased self-assessments and, correspondingly, biases in their perceptions of how their peers view them. Previous research has documented such gender biases in undergraduate science students’ self-assessments. One study, for example, found that women who earn high grades in their physics class report similar levels of self-efficacy (i.e., beliefs about their capabilities to successfully carry out physics tasks) as men who earn much lower grades [32]. Another study observed that men predict higher chemistry exam grades than women despite earning similar exam grades [33].
@@
These and similar lines of research will further illuminate how students internalize their peer recognition and why men and women internalize their peer recognition differently, as we observe in the large-scale analysis presented here. With these mechanisms, the research community can then design, implement, and evaluate instructional interventions that are effective in mitigating gender differences in perceived peer recognition. Such interventions will help more women and students from marginalized groups develop their physics identity and better succeed in their physics courses.